OpenStudy (starlight513):

The average score on a nationwide standardized test is found to be 124 with a standard deviation of σ = 8.4. We at GCA believe our students scored better than average on the test. A sample of 40 GCA student's scores was taken. The average score for the sample was 126, A hypothesis test is conducted at the .05 significance level. Which of the following is the correct test statistic for the hypothesis test? z = -1.51 z = 1.51 z = 2.65 z = -2.65

Directrix (directrix):

z = (x-bar - µ)/(σ/sqrt(n)) z = (126 - 124)/(8.4/sqrt(40)) z = 1.51